OpenStudy (anonymous):

Calculus:

OpenStudy (anonymous):

What is: \[\int\limits_{?}^{?}xe ^{-x ^{2}}dx\]

OpenStudy (anonymous):

thank you. can u please explain how you go that?

OpenStudy (blurbendy):

yes, first use u-substitution so you let u = -x^2 du = -2xdx xdx = -1/2du -1/2 (integral sign) e^u du = - (e^u) / 2 + C = - (e^-x^2) / 2 + C You can simplify that to: - 1 / (2ex^2) + C Understand?

OpenStudy (anonymous):

yes! thank you!

OpenStudy (blurbendy):

no problem!